### v2

# QSTN: A Modular Framework for Robust Questionnaire Inference with Large Language Models

Abstract
We introduce QSTN, an open-source Python framework for systematically generating responses from questionnaire-style prompts to support in-silico surveys and annotation tasks with large language models (LLMs). QSTN enables robust evaluation of questionnaire presentation, prompt perturbations, and response generation methods. Our extensive evaluation (>40 million survey responses) shows that question structure and response generation methods have a significant impact on the alignment of generated survey responses with human answers. We also find that answers can be obtained for a fraction of the compute cost, by changing the presentation method. In addition, we offer a no-code user interface that allows researchers to set up robust experiments with LLMs without coding knowledge. We hope that QSTN will support the reproducibility and reliability of LLM-based research in the future.
QSTN: A Modular Framework for Robust Questionnaire Inference
with Large Language Models
Maximilian Kreutner1, Jens Rupprecht1, Georg Ahnert1, Ahmed Salem1, Markus Strohmaier1,2,3 1University of Mannheim, 2GESIS - Leibniz Institute for the Social Sciences, 3CSH Vienna
1 Introduction
Figure 1: QSTN Facilitates Easy To Customize and Robust Questionnaire Inference with LLMs. QSTN provides a fully modular pipeline with different ways to present the questionnaire, prompt perturbations and to choose a response generation method, with automatic parsing. Both local and remote inference are supported.
Questionnaires have become an important format to probe, assess, and utilize large language models (LLMs) via prompts. Questionnaire-like prompts have been a popular way to evaluate LLMs on tasks such as common knowledge understanding Hendrycks et al. (2021), language comprehension Hu et al. (2023); Sravanthi et al. (2024); Kim et al. (2024), and mathematical reasoning Satpute et al. (2024); Wei et al. (2023). Other work uses existing questionnaires to evaluate LLMs’ values; for example, political bias Röttger et al. (2024); Rozado (2024), personality traits Jiang et al. (2024); Shu et al. (2024); Pellert et al. (2024), or psychometric profiles Ye et al. (2025). With the increasing capability of LLMs, researchers have found additional use cases, such as the creation of synthetic survey responses Argyle et al. (2023); Ma et al. (2024) or data annotation Tan et al. (2024).
Despite the widespread use of questionnaire-like prompts, concerns have been raised about the robustness of LLM responses to such prompts. The closed-ended responses of an LLM can vary strongly from its open-ended responses Röttger et al. (2024); Wang et al. (2024), LLM responses can be biased towards specific survey response options Tjuatja et al. (2024); Rupprecht et al. (2025), and downstream performance is strongly affected by small changes in the questionnaire configuration Cummins (2025); Ahnert et al. (2025).
To address and investigate some of these concerns, we present QSTN (pronounced “Question”) - a Python framework designed to facilitate the execution of questionnaire-style experiments with LLMs. QSTN simplifies the process of creating robust variations of question prompts and answer generation methods, thereby facilitating reproducibility and the analysis of the reliability of LLM-based questionnaire research. QSTN provides a complete, modular pipeline, as depicted in Figure 1, for creating the questionnaire presentation, adjusting various parts of the prompt with perturbations, choosing the response generation method, performing inference, and finally, parsing the generated text. We evaluate our framework on more than 40 million survey responses and find that the controlled variation of the experiment pipeline can increase the alignment of generated responses with human survey answers and that the responses can be obtained for a fraction of the compute cost.
2 Core Features
QSTN was developed with three objectives in mind: First, it enables robust evaluation of and with LLMs, addressing prompt sensitivity Tjuatja et al. (2024); Dominguez-Olmedo et al. (2024). QSTN is engineered to address this challenge directly through a highly modular and configurable design. Each part of the pipeline can be exchanged independently from the other parts.
Second, QSTN is designed to be efficient, so it can be used in large-scale studies. For experiments with multiple prompt variations and/or personas, we automatically utilize prefix caching and batching for local inference in vLLM Kwon et al. (2023), and asynchronous calling with the AsyncOpenAI API OpenAI (2023).
Finally, QSTN is designed to be as easy to use as possible. Since we maintain the common prompt format of the system prompt and user prompt, adapting a project to QSTN is seamless. The package offers a complete pipeline from prompt creation and inference to parsing, which can be done in only three function calls to the package. Integration with existing vLLM and OpenAI packages is straightforward.
QSTN’s core strength lies in its ability to systematically and easily control and vary the setup of questionnaire-like prompting experiments. The following aspects of the experiments can be exchanged and varied by simply switching out one module for another.
2.1 Questionnaire Presentation
QSTN supports three distinct questionnaire presentation modes, as shown in Figure 2:
  • Sequential: Each question is asked in the same conversation context in multiple, sequential chat calls.
  • Battery: All questions are asked in   battery and the model is expected to answer all questions in one response in the same context.
  • Single-item: Each question is asked in a new context, with the LLM not being aware of the previous questions and answers.
Figure 2: QSTN Questionnaire Presentation Modes
Questionnaire presentation is a fundamental decision to make when using LLMs with questionnaire-like prompts. For example, if we want to annotate data, is it better to give all annotation questions in the same prompt, or should each question be asked in a new context? There is evidence that keeping multiple tasks in the prompts can improve variety for creative writing Zhang et al. (2025) and improve performance for classification tasks in moral foundations Chen et al. (2025). LLMs are also able to perform multiple tasks of different kinds in   battery Son et al. (2024), which can save computing time.
2.2 Prompt Perturbation
Previous studies found that LLMs synthetic survey responses are highly sensitive to prompt perturbations and exhibit biases, such as token biases, recency bias, or A-bias Pezeshkpour and Hruschka (2024); Li and Gao (2025); Rupprecht et al. (2025); Dominguez-Olmedo et al. (2024); Röttger et al. (2024). QSTN can automatically randomize or reverse both the order of the questions within the survey and the order of answer options for each question to identify and mitigate these biases. This ensures that high performance is robust and independent of ordering. Previous research has found that LLMs can be sensitive to small changes in prompt format He et al. (2024); Sclar et al. (2024). QSTN allows users to define custom answer label schemas (e.g., A/B/C, 1/2/3, i/ii/iii), enabling rigorous testing of a model’s robustness to superficial formatting changes. QSTN can perform the following Answer Option Perturbations, which are shown in Figure 3:
Figure 3: QSTN Supported Prompt Perturbations
  • Reversed Response Order: The order of answer options is reversed (e.g., a scale from ‘1: Very important’ to ‘5: Not important’ becomes ‘1: Not important’ to ‘5: Very important’).
  • Missing Refusal Option: The “Don’t know” or refusal option is removed from the list of choices.
  • Odd/Even Scale Transformation: For scales with an even number of options, a semantically appropriate middle category is added, transforming it into an odd-numbered scale (e.g., by adding ‘Neutral’). Conversely, for odd-numbered scales, we remove the middle category to create an even scale and adjust the integer label accordingly.
In addition, QSTN can perform the following Question Perturbations:
  • Typographical Errors: three types of typos can be introduced: Key Typo (replacing a character with a random one), Letter Swap (swapping two adjacent characters in a random word), and Keyboard Typo (replacing a character with an adjacent one on a QWERTY keyboard).
  • Semantic Variations: Additional semantic variations can be introduced while preserving the original meaning: first, by Synonym Replacement, where a variable amount of words in the original question are replaced with synonyms. Second, through Paraphrasing the entire question is rephrased.
2.3 Response Generation
While generative language models are designed to generate open-ended text, previous studies have implemented various approaches to constrain LLMs to closed-ended responses (e.g., Ma et al., 2024). We define Response Generation Methods as techniques used to elicit closed-ended responses from large language models to questionnaires Ahnert et al. (2025). QSTN supports the following Response Generation Methods, with examples being shown in Figure 4:
  • Token Probability-Based Methods: Extract probabilities for response options from the output token probabilities of an LLM.
  • Restricted Generation Methods: Force the model to respond only with designated response options using formatting instructions in the prompt and (optionally) restrict the vocabulary of the LLM through structured outputs.
  • Open Generation Methods: Generate open-ended responses first and then classify them in a second step.
Figure 4: QSTN Response Generation Methods
The Restricted Generation Methods can be used to generate exactly one of the available response options—optionally in JSON format, or with reasoning—or to generate a verbalized distribution of probabilities for all response options, following Meister et al. (2025). All Response Generation Methods can be adjusted to, e.g., have the model generate a prefix before token probabilities are extracted. QSTN includes suitable parsers for all generated responses: JSON & LLM-as-a-judge.
3 Using QSTN
The package containing QSTN can be installed in the desired environment using pip. We support both a lightweight installation with pip install qstn, which only requires dependencies for API usage, and a full installation with pip install qstn[vllm] , which allows for local inference with vllm. QSTN is easily integrable into current workflows, requiring just a total of three function calls for the most basic functionality, and it still allows users to freely define their prompts. A minimum usage example is given in Listing 5. By simply exchanging the function in the inference step, the questionnaire presentation can be adjusted, or a different type of parser can be selected. Additionally, building on this simple example, only one more module is needed to implement controlled prompt perturbations and response generation methods.
import qstn
import pandas as pd
from vllm import LLM
# 1. Prepare questionnaire and persona data
questionnaires = pd.read_csv("hf://datasets/qstn/ex/q.csv")
personas = pd.read_csv("hf://datasets/qstn/ex/p.csv")
prompt = (
[…]
f"{qstn.utilities.placeholder.PROMPT_QUESTIONS}"
interviews = [
qstn.prompt_builder.LLMPrompt(
questionnaire_source=questionnaires,
system_prompt=persona,
prompt=prompt,
) for persona in personas.system_prompt]
# 2. Run Inference
model = LLM("Qwen/Qwen3-4B", max_model_len=5000)
results = qstn.survey_manager.conduct_survey_single_item(
model, interviews, max_tokens=500
# 3. Parse Results
parsed_results = qstn.parser.raw_responses(results)
Figure 5: Minimum usage example of QSTN. QSTN can be easily integrated into existing projects, requiring just three function calls to operate. Users familiar with vllm or the OpenAI API can use the same Model/Client calls and arguments. In this example reasoning and the generated response are automatically parsed.
Non-Code User Interface
QSTN offers a User Interface to create and run inference with LLMs without having to program any Python code. The UI offers the same core functionality as the main framework, allowing users to upload questionnaires, systematically alter the prompt structure, set model parameters, and run inference. While the UI generally offers the same functions as the coding package, some more advanced features, such as inferencing models directly through the vllm API, are currently not supported.
4 Evaluation
We evaluate QSTN primarily on the task of generating synthetic survey responses, which is a topic of growing interest. Our results demonstrate that our proposed variations significantly influence both the alignment of synthetic data with real-world responses and computational efficiency. Across all experiments, we use the following instruction-finetuned versions of the models: LLama3 1-70B Grattafiori et al. (2024), Qwen3 4-30B Yang et al. (2025), Phi-4-mini Abdin et al. (2024), Gemma3 4-27B Team et al. (2025), OLMo2 1-32B OLMo et al. (2024), Yi1.5 6B Young et al. (2024), and Gemini1.5 Pro Team et al. (2024). We present new results and evaluations regarding questionnaire presentation and provide an overview of previous experiments that were conducted in or implemented into QSTN, which evaluate Prompt Perturbations and Response Generation Methods.
4.1  Questionnaire Presentation
We start by demonstrating that the presentation of the questionnaire significantly impacts the subpopulation alignment of generated responses with real answers. Furthermore, selecting the optimal method results in savings for both token usage and GPU time. We test three fundamentally different presentations, as described in Section 2.1.
We base our experiments on Bisbee et al. (2024), where respondents of the ANES survey are instructed to consider a certain group and to indicate the degree to which they experience warm (positive, affectionate, etc.) or cool (negative, disdainful, etc.) feelings toward members of that group on a scale from 0 to 100. For each of the 7530 participants, we use three different seeds, which leads to a total of 10,843,200 individual question responses across 16 questions, 10 different models, and 3 different presentations.
We use the same prompts as in the initial study, with the addition of an instruction on how to format the output to align with the response generation method. Our full prompts can be seen in Table 7 in the Appendix. Respondents were stratified into subpopulations based on the intersection of gender, race, and ideology (see Appendix Table 6 for full subpopulation attributes). We measure individual alignment via Mean Absolute Error and subpopulation distributional alignment via Wasserstein distance; results are displayed in Table 1. To quantify the effects of questionnaire presentation, we fitted Ordinary Least Squares and Weighted Least Squares models for MAE and Wasserstein distance, respectively. Both models include interaction terms between presentation and model, as well as fixed effects for iteration seeds. The   single-item presentation and Llama-3.3-70B-Instruct serve as the reference categories.
Mean Absolute Error $\downarrow\downarrow$ Wasserstein distance $\downarrow\downarrow$ questionnaire presentation sequential battery single-item sequential battery single-item gemma-3-4b-it 20.96 0.02 21.92 0.01 19.94 0.02 16.48 0.01 17.62 0.02 16.39 0.01 gemma-3-12b-it 18.26 0.01 18.07 0.02 19.11 0.01 14.53 0.02 13.44 0.01 16.44 0.01 gemma-3-27b-it 17.59 0.01 17.90 0.01 18.01 0.01 14.00 0.01 14.26 0.01 15.17 0.00 Llama-3.2-1B-Instruct 30.89 0.25 30.22 0.07 35.69 0.12 18.66 0.33 18.15 0.12 27.52 0.17 Llama-3.2-3B-Instruct 24.20 0.10 22.98 0.04 24.32 0.06 13.14 0.11 13.50 0.03 15.88 0.07 Llama-3.1-8B-Instruct 21.01 0.04 20.88 0.02 20.87 0.04 13.62 0.02 12.90 0.02 14.11 0.04 Llama-3.3-70B-Instruct 18.23 0.00 17.67 0.00 17.87 0.01 14.18 0.00 13.56 0.01 14.73 0.01 Phi-4-mini-instruct 20.98 0.03 19.72 0.01 21.23 0.03 11.69 0.06 12.21 0.02 14.56 0.05 Qwen3-4B-Instruct-2507 19.29 0.02 20.34 0.01 20.05 0.02 13.75 0.02 15.59 0.01 15.18 0.01 Qwen3-30B-A3B-Instruct-2507 17.68 0.02 17.67 0.01 18.29 0.01 13.88 0.02 13.68 0.01 15.21 0.02
Table 1: Individual and subpopulation alignment based on   questionnaire presentation. Mean absolute error for each individual response and weighted mean Wasserstein distance across the subpopulations. Wasserstein distance significantly improves with sequential and battery presentation for most models, compared to single-item.
We find that questionnaire presentation has a substantial impact on distributional alignment, whereas the effects on individual-level accuracy, while statistically significant, are practically marginal. For the reference model, the   battery presentation yields the strongest improvement in subpopulation alignment ($\beta_{\text{WD}}=-1.17,p<0.01$), representing an approximate $8\%$ better alignment than with the   single-item presentation. This effect is consistent across the large models we tested, as the interaction effect for both Qwen-30B and Gemma-27B was not statistically significant. However, for smaller models, the effect is highly architecture-dependent: Phi-4-mini achieves the best overall alignment in our experiment using the   sequential presentation, whereas gemma-3-4b achieves the best alignment with   single-item presentation.
Considering the large differences in tokens and compute time between the presentation methods (shown in Table 2), we recommend the   battery presentation as the default for future questionnaire-based experiments with large persona prompts. However, thorough tests should be conducted to ensure that performance is comparable to other presentations for the specific model and task at hand. QSTN makes these validation experiments accessible by requiring just a single method change in the pipeline.
4.2  Prompt Perturbation
In previous research Rupprecht et al. (2025), we found a consistent recency bias in all nine models tested, favoring the same answer option when placed at the end of the options list instead of the beginning. This effect was substantial, with the selection frequency of the semantically same option increasing by more than 20 times for Llama-3.1-8B when moved to the last position, while all other configurations, such as question and prompt phrasing, were kept constant.
All models facing prompt perturbations showed some level of non-robust responses, whereas larger models such as Llama-3.3-70B and Gemini-1.5-Pro respond more robustly. The magnitude of the effect of perturbations (e.g., on the answer option or the question phrasing) on response robustness mainly depends on the type of perturbation applied. We identified that some of the Answer Option Perturbations and Question Perturbations have a larger impact on response robustness than others (see Table 3). Reversing the answer options or introducing typos or paraphrasing the questions is more harmful to robustness than swapping characters within a word or removing the refusal category. In addition, we found that 67% and 89% of models select the middle category significantly more often when a 5- or 11-point Likert scale is provided, respectively.
These findings underline the importance of robustness checks, e.g., through prompt perturbations. QSTN allows the user to apply various perturbations automatically to any questionnaire presented and thus assess the response robustness of the LLM.
Presentation Calls Input T. Output T. Inference Time   sequential 16 8216 288 09:29:05   battery 1 723 142 01:34:45   single-item 16 4288 288 03:22:23
Table 2: API Calls, Tokens and inference time of different   questionnaire presentations. We report the number of API calls, tokens and inference time for the largest model LLama-3.3-70B-Instruct. The tokens are calculated on one persona and the time is measured by a whole run of 7530 personas with 3 seeds. All experiments have been conducted with vllm on two 2 NVIDIA H100 GPUs (tensor-parallel).
4.3  Response Generation Methods
To investigate the impact of Response Generation Methods on generated questionnaire responses, we predict survey responses to questions of political attitudes in the American National Election Study ANES (2016), the German Longitudinal Election Study GLES (2017, 2025), and the American Trends Panel ATP (2021). We thereby partially replicate the studies by Argyle et al. (2023), von der Heyde et al. (2025), and Santurkar et al. (2023), while extending them to include additional Response Generation Methods. We compare 8 Response Generation Methods on 10 open-weight LLMs, including reasoning models. For robustness, we include 4 prompt variations, 3 random seeds for temperature-scaled decoding, as well as greedy decoding. Overall, we simulate 32 mio. survey responses with QSTN, and evaluate their alignment with human survey responses on individual and subpopulations levels. For subpopulation-level alignment, we split the set of respondents into subpopulations by considering all unique values of all persona attributes that were included in the studies we replicate, e.g., women & men, people from different states, etc. We report the subpopulation-level alignment on categorical response distributions using total variation distance (see also Meister et al., 2025; Baan et al., 2022).
 Answer Options  Question Perturbations
Model (1) (2) (3) (4) (5) (6) (7) (8)
Llama-3.3-70B 0.50 0.73 0.60 0.52 0.76 0.58 0.58 0.66
Llama-3.1-8B 0.08 0.39 0.27 0.32 0.31 0.23 0.32 0.16
Llama-3.2-3B 0.10 0.11 0.16 0.10 0.16 0.18 0.23 0.10
Llama-3.2-1B 0.00 0.11 0.03 0.05 0.11 0.00 0.13 0.02
Gemini-1.5-Pro 0.69 0.76 0.55 0.68 0.73 0.66 0.60 0.55
Phi-3.5-mini 0.53 0.81 0.45 0.50 0.61 0.47 0.71 0.53
Mistral-7B-v0.3 0.68 0.81 0.53 0.58 0.65 0.60 0.71 0.53
Qwen-2.5-7B 0.32 0.48 0.45 0.48 0.65 0.45 0.55 0.44
Yi-1.5-6B 0.47 0.68 0.55 0.50 0.50 0.45 0.65 0.29
Table 3: Impact of  Answer Option and  Question Perturbations on the Response Robustness of different LLMs ($\uparrow$)($\uparrow$). Share of fully robust responses per model. Bold indicates the highest robustness score for that perturbation type. Perturbation Keys: (1) Reversed Answer Options, (2) Missing Refusal, (3) Even Scale, (4) Key Typos, (5) Letter Swap, (6) Keyboard Typos, (7) Synonyms, (8) Paraphrase
Table 4 shows selected OLS regression coefficients for subpopulation-level alignment. We find that the Verbalized Distribution Method yields significant improvements on most datasets. In combination with the individual-level alignment results presented in Ahnert et al. (2025), we conclude that: (i) the choice of Survey Response Generation Method should be well-justified for in-silico surveys, since we find significant differences between these methods. (ii) We do not recommend the use of Token Probability-Based Methods, as they generate misaligned survey responses. (iii) For predicting closed-ended survey responses, we suggest to consider Restricted Generation Methods first, as they consistently show significant improvement over other methods while also being more computationally efficient than Open Generation Methods.
Response ANES GLES GLES ATP
Generation Method 2016 2017 2025 2021
Intercept .374* .312* .288* .503*
  First-Token Prob. -.003 .147* .194* -.049*
  Verbalized Distrib. -.074* -.057* -.013 -.168*
  Open-Ended Distrib. -.006 -.052* -.037* -.082*
Table 4: Impact of  Response Generation Methods on Subpopulation-Level Alignment ($\downarrow$)($\downarrow$). OLS regression coefficients by dataset with total variation distance ($\downarrow$)($\downarrow$) as the dependent variable and Survey Response Generation Method, prompt perturbation, and LLM as independent variables. We show coefficients for selected Response Generation Methods (Reference: Restricted Choice)—see Appendix B for all coefficients and more details on OLS model choice. The Verbalized Distribution Method leads to significant improvements. $\text{*}\,p<0.05\;\text{(Benjamini–Hochberg corrected)}$
5 Related Work
Due to the importance of controlled prompt perturbation, a number of frameworks have started to address this issue. In general, QSTN supports controlled variation and combines it with the pipeline to allow for automatic parsing of all prompt variations. Additionally, as QSTN allows for modular prompts, these frameworks can be used in conjunction with it. PromptSuite Habba et al. (2025) focuses on prompt perturbation through paraphrasing and formatting. PromptSource Bach et al. (2022) is a framework for making and sharing different types of natural language prompts. Prompt-Agnostic Fine-Tuning (PAFT) Wei et al. (2025) varies prompts in the fine-tuning process rather than during inference.
There are also frameworks that model the entire pipeline of LLM experiments, similar to QSTN. Unitxt Bandel et al. (2024) is an open-source Python framework for data processing pipelines. While powerful, it requires users to understand the Unitxt operator language, which can add cognitive overhead. The EDSL framework Horton and Horton (2024) can be used to run surveys with LLMs, but it does not provide full freedom over the exact system prompt or prompt and the Response Generation Method.
6 Conclusion
We introduce QSTN, a Python framework designed to make LLM inference with questionnaires more robust. Our evaluation demonstrates that by enabling controlled variations in the generation process, QSTN can significantly improve the alignment of generated responses with human answers while reducing inference costs. A core feature of QSTN is its modularity, allowing researchers to easily vary their experimental setup with only minimal additional coding effort. The framework is broadly applicable to tasks such as data annotation, synthetic data generation, persona studies, and the analysis of LLM behavior itself.
Limitations
Currently, our evaluation is primarily focused on the creation of synthetic survey responses. We hope that by releasing QSTN to the open-source community, more robust experiments can be conducted in other application domains. While we support a variety of different Response Generation Methods and parsing options, we currently do not support every type of structured output; for example, we do not support output that is guided by a regex pattern or context free grammar. As such, not every type of experiment can currently be conducted in QSTN. We hope that by making the project open-source, we will be able to support more ways to conduct experiments. Additionally, while we plan to add support for non-instruct models, they are currently not supported.
Appendix A Questionnaire Presentation
As another measure of individual evaluation of predictions, we calculate the Pearson correlation between predictions and the ground truth and present it in Table 5. Similarly to the Mean Absolute Error, we see little difference in the performance of the different questionnaire presentations.
questionnaire presentation sequential battery single-item gemma-3-4b-it 0.59 ± 0.00 0.55 ± 0.00 0.57 ± 0.00 gemma-3-12b-it 0.62 ± 0.00 0.62 ± 0.00 0.64 ± 0.00 gemma-3-27b-it 0.62 ± 0.00 0.61 ± 0.00 0.61 ± 0.00 Llama-3.2-1B-Instruct 0.25 ± 0.01 0.18 ± 0.00 0.10 ± 0.00 Llama-3.2-3B-Instruct 0.51 ± 0.00 0.49 ± 0.00 0.52 ± 0.00 Llama-3.1-8B-Instruct 0.56 ± 0.00 0.57 ± 0.00 0.56 ± 0.00 Llama-3.3-70B-Instruct 0.64 ± 0.00 0.64 ± 0.00 0.64 ± 0.00 Phi-4-mini-instruct 0.48 ± 0.00 0.49 ± 0.00 0.52 ± 0.00 Qwen3-4B-Instruct-2507 0.60 ± 0.00 0.55 ± 0.00 0.59 ± 0.00 Qwen3-30B-A3B-Instruct-2507 0.62 ± 0.00 0.62 ± 0.00 0.59 ± 0.00
Table 5: Mean and Standard Deviation of Pearson Correlation between Prediction and Ground Truth. Similar to Mean Absolute Error, individual alignment measured with Pearson Correlation shows little difference between different questionnaire presentations.
We show all attributes we considered for the subpopulation analysis for Wasserstein distance in Table 6. The full regression results for both MAE and Wasserstein Distance can be seen in 8. We report the coefficients and the Benjamini-Hochberg corrected p-values. Additionally, we want to determine if the questionnaire presentation has different effects on different questions. For this, we fit an additional Weighted Least Squares regression on all subpopulations based on the full interaction between the questionnaire presentation, the model, and the specific interview question. We set   single-item, the biggest model Llama-3.3-70B-Instruct and the first question as the reference categories, as for this question the LLM has no answers for the other questions in context regardless of the questionnaire presentation.
All questions show improvements, and a subset of five questions shows statistically significant improvement $(p<0.05)$ when using   battery presentation instead of   single-item presentation. The largest improvement is in the question about feelings towards the group of Gays and Lesbians $(\beta=-3.82,p<0.01)$ when using   battery presentation. Figure 6 visually confirms this: when previous questions and answers are included in the context, the model’s response distribution aligns much more closely with the ground truth, exhibiting a similar tendency toward neutral answers. The other significant questions concern the groups of White Americans, Asian Americans, Christians, and Liberals.
Figure 6: Answer Distributions. Predictions and Ground Truth Distributions across the whole population compared for LLama-3.3-70B-Instruct and the question “How do you feel towards Gays and Lesbians?”. We can see a clear shift towards the middle for this question, when models are given context of the previous questions and answers, which aligns more closely with human answers.
Attribute Category
Race Non-Hispanic White
Non-Hispanic Black
Hispanic
Gender Male
Female
Ideology Extremely Liberal
Liberal
Slightly Liberal
Moderate
Slightly Conservative
Conservative
Extremely Conservative
Table 6: Subpopulations: We consider these subpopulations for analysis. We have the same subpopulations as the initial study by Bisbee et al. (2024).
We use the same prompt as that used in Bisbee et al. (2024), as displayed in Table 7. We adjust the output instructions to fit our choice response generation method and add all questions as instructions in the   battery presentation. For all models we use the default hyperparameter settings.
[…]
Table 7: Prompt. We use the same prompts for   sequential and   single-item and a slightly modified output instruction for the   battery presentation. For   battery presentation we ask all questions separated by new lines.
(1) (2)
MAE (OLS) WD Score (WLS)
Questionnaire
Presentation   sequential 0.362** -0.546*
  battery -0.199** -1.166**
Model Llama 3.1 8B 2.999** -0.617*
Llama 3.2 1B 17.822** 12.796**
Llama 3.2 3B 6.450** 1.152**
Phi-4 Mini 3.366** -0.163
Qwen3 30B 0.420** 0.488
Qwen3 4B 2.187** 0.454
Gemma 3 12B 1.245** 1.713**
Gemma 3 27B 0.142** 0.448
Gemma 3 4B 2.074** 1.662**
Interactions
(Presentation Model)   sequential Llama 3.1 8B -0.216** 0.060
  battery Llama 3.1 8B 0.213** -0.044
  sequential Llama 3.2 1B -5.162** -8.311**
  battery Llama 3.2 1B -5.273** -8.203**
  sequential Llama 3.2 3B -0.476** -2.195**
  battery Llama 3.2 3B -1.136** -1.211**
  sequential Phi-4 Mini -0.619** -2.322**
  battery Phi-4 Mini -1.318** -1.183**
  sequential Qwen3 30B -0.973** -0.784*
  battery Qwen3 30B -0.418** -0.372
  sequential Qwen3 4B -1.125** -0.885*
  battery Qwen3 4B 0.490** 1.574**
  sequential Gemma 3 12B -1.218** -1.366**
  battery Gemma 3 12B -0.843** -1.832**
  sequential Gemma 3 27B -0.779** -0.625
  battery Gemma 3 27B 0.087 0.250
  sequential Gemma 3 4B 0.652** 0.637
  battery Gemma 3 4B 2.175** 2.394**
Table 8: Regression Results for MAE and Wasserstein Distance. ($\downarrow\downarrow$) Model (1) uses OLS on Mean Absolute Error. Model (2) uses WLS on Wasserstein Distance, weighted by subpopulation count. Significance levels are based on Benjamini–Hochberg corrected p-values. We can see significant effects for both the questionnaire presentation, but also for the interaction between smaller models and the presentation. Reference categories: Presentation:   single-item, Model: Llama-3.3-70B-Instruct. $\text{*}\,p<0.05,\ \text{**}\,p<0.01$
Appendix B Response Generation OLS Regressions
We obtain the subpopulation-level alignment for each simulation specification and subpopulation, as described in Section 4.3. To identify significant differences in survey response alignment between the response generation methods, we fit the following OLS regression model separately on each dataset (see Table 9): We use the per-subpopulation total variation distance ($\downarrow$)($\downarrow$) as the dependent variable and Survey Response Generation Method (reference: Restricted Choice), LLM (reference: Llama 8B), and prompt perturbation (reference: Full Text response options) as independent variables. We use cluster-robust SEs, clustering by seed × decoding strategy, which allows for arbitrary correlation and heteroskedasticity within clusters while assuming independence across clusters. This appropriately reflects the repeated-measures structure of our evaluation. We do not include interaction terms into the OLS model to mitigate multicollinearity—all VIF values are < 3. We apply Benjamini–Hochberg correction across all reported coefficients in all datasets. Key coefficients for the Verbalized Distribution Method, as well as OLMo 32B and Qwen 32B remain significant even under Bonferroni correction, although Bonferroni is known to be overly conservative in regression settings with correlated predictors.
ANES 2016 GLES 2017 GLES 2025 ATP 2021
Intercept 0.374** 0.312** 0.288** 0.503**
Response
Generation Method   First-Token Probabilities -0.003 0.147** 0.194** -0.049*
  First-Token Restricted 0.064** 0.220** 0.234** -0.005
  Answer Prefix -0.002 0.047* 0.085** -0.082**
  Restricted Reasoning 0.017 -0.035* -0.026 -0.084**
  Verbalized Distribution -0.074** -0.057** -0.013 -0.168**
  Open-Ended Classif. 0.026 -0.011 -0.027 -0.051**
  Open-Ended Distrib. -0.006 -0.052** -0.037* -0.082**
Model Llama 3B -0.051* 0.031 0.066** -0.039*
Llama 70B -0.052* -0.089** -0.127** 0.007
OLMo 1B -0.023 0.109** 0.114** 0.109**
OLMo 7B -0.062** 0.070** 0.077** -0.030
OLMo 32B -0.070** -0.073** -0.109** 0.016
Qwen 8B 0.016 0.020 -0.050* 0.075**
Qwen 8B with Reasoning -0.012 0.002 -0.010 0.019
Qwen 32B -0.076** -0.108** -0.161** -0.036*
Qwen 32B with Reasoning -0.056** -0.067** -0.081* -0.106**
Response Option
Variants Full Text, Reversed 0.001 -0.005 0.037 -0.003
Indexed 0.010 0.003 0.000 0.022*
Indexed, Reversed 0.035* 0.011 0.026 0.030**
Table 9: Impact of  Response Generation Methods on Subpopulation-Level Alignment ($\downarrow$)($\downarrow$). OLS regression coefficients by dataset with total variation distance ($\downarrow$)($\downarrow$) as the dependent variable and Survey Response Generation Method (reference: Restricted Choice), LLM (reference: Llama 8B), and prompt perturbation (reference: Full Text response options) as independent variables. The Verbalized Distribution Method and larger models lead to significant improvements. $\text{*}\,p<0.05,\ \text{**}\,p<0.01\;\text{(Benjamini–Hochberg corrected)}$